OpenStudy (jokersmart):

Check my answer? Choose all that apply. Dennis is entering a contest to guess the number of marbles inside a cylinder. He wants to find the volume of the marbles and the volume of the cylinder to estimate how many marbles might be in the cylinder. Which formulas does he need to make a good guess? Choose all answers that are correct. (In these formulas, r represents radius, h represents height, l represents length, and w represents width.) A) V=lwh B) V= 4/3 pi r^3 <-- C) V= pi r^2 h <--- D) V= 1/3 pi r^2 h

OpenStudy (jokersmart):

\[A. V=lwh\] \[B. V=\frac{ 4 }{ 3 }\pi r ^{3}\] \[C. V=\pi r ^{2}h\] \[D. V=\frac{ 1 }{ 3 }\pi r ^{2} h\]